28.) Two charged pellets, each with a charge of 5C, are located a distance 15m apart. What is the electric force between the two pellets?
Answer #1

Coulomb's law is a physical law that describes the force of interaction between fixed point electric charges depending on the distance between them.

F = K*Q1*Q2/r²

where K = 8.98 * 10^9 N.m²/C²

F = 8.98 * 10^9 * 5 * 5/15²

F = 0.998 * 10^9 ~ 10^9 N

Thus the magnitude of the electric force between two pellets is 10^9 N.